titanium metal from its oxide according to the following balanced equation: TiO2 + 2C >> Ti(s)+ 2CO(g) when 30.6 kg of C reacts with 76.2 kg of TiO2, 39.8 kg of Ti is produced. Find the limiting reactant, theoretical yield ( in kg) and the percent yield.
1
Expert's answer
2019-04-17T04:16:21-0400

"TiO2(s)+2C(s) \to 2CO(g)+Ti(s)"

"n(C)=\dfrac{30600g}{12g/mol}=2550moles"

"n(TiO2)=\dfrac{76200g}{79.86g/mol}=954.17moles(limiting)"

"\dfrac{76200g*47.867g/mol}{79.86g/mol }=45673.25g"

"%yield=\dfrac{39800g}{45673.25g}*100=87.14%"

"%yield=" %yield="\dfrac{39800g}{45673.25}*100=87.14" %